31.3 L of concentrated nitric acid (8.50 mol L-1) was spilled on a road. Sodium carbonate solid was spread on the acid to react as in the equation:

CO32-(aq) + 2H+ → CO2(g) + H2O

How many grams of sodium carbonate are needed to neutralise the acid?


1
Expert's answer
2022-03-24T22:54:04-0400

If 1litre= 8.50mol

31.3×8.50= 266.05moles

Mole ratio= 1:2


Moles of sodium carbonate= (1/3)×266.05

= 88.68moles

Mass= 106× 88.68= 9400.43g